To: J. Albertine, D. Helms, W. Skinner
cc: Division (M7), FEL Coordination Group
From: F. Dylla
Subject: IRFEL Weekly Report, August 18-22, 1997
Date: August 22, 1997
Highlights for the week include successful measurement and acceptance
of the optical chiacane dipole magnets and completion and pumpdown
of the gun beamline hardware to the injector cryounit.
Management attention was focused on the completion of the financial
reports for the July Monthly Report, which were distributed to
DOE and the Navy on Friday (Aug.22), and additionally on cost/schedule
planning of the FY98 commissioning and upgrade activities.
Congressman Bateman's Office requested an update on the current
status of the FEL project. A telephone briefing to the Congressman's
staff was followed by a brief written summary of the conversation.
Revisions of pre-proposals that are being prepared for submission
to the DOE-BES Program Office were received this week from collaborators
at NCSU, College of William and Mary, and the University of Michigan.
Discussions were held with the Administration of Old Dominion
University on the fit-up of the ODU Laboratories in the ARC Building.
A draft document is being prepared that will deal with the safety
and environmental issues for the university tenants' ARC laboratories.
In addition, we discussed co-locating the Jefferson Laboratory
Optics Laboratory with ODU's planned Laser Laboratory.
Regarding the beam-transport system:
Due to the brief report on beam transport hardware last week.
we have included an expanded report this week.
The optical chicane dipoles (DW) had their bolts re-tightened
and had some readjustment of the field clamp position. With these
relatively minor actions, all magnets meet specification for core
field, K1 value and field integral through the centerline. The
first of these adjusted magnets tested on the moving coil apparatus,
easily met the one part in ten thousand uniformity specification
for field integral across the 9 cm required good field region.
The design is thereby validated.
Two DW magnets are exceptions in their mechanical measurements,
having local variation of the mu metal glue-down/gap at the .002
inch level. The one tested on the moving coil apparatus had 3
parts in ten thousand field variation in the field integral across
the 9 cm required good field region at most higher excitation
levels. We will be working to improve these two magnets. We
are keeping Everson Electric, the assembly vendor, informed of
all these minor problems.
The next set of 9 dipoles slated for assembly is the Injection/Extraction DU/DV styles. After a one week delay due to lack of materials due to the UPS strike, Everson is gluing mu metal at a one magnet per day rate starting last week and has one DU and one DV assembled up to the point of a back-ordered terminal block. A cache of 38 of the blocks was located and sent by overnight mail to Everson. The core slabs, back legs and parts for the 180 degree dipoles are complete and are being pinned together. They are slated for delivery next week. Winding of the last of four 180 degree dipole coils started again as the vendor received a reel of copper that had inadvertently not been sent with the initial shipment. Potting will follow, with delivery of a first set of coils possible in two weeks. Details on the remaining parts of the magnet measurement apparatus for these magnets reached the sign-off stage.
In the quadrupole and sextupole area, winding of coils for the
eight Panofsky trim quads started at Magnet Enterprises International
(MEI). Mechanical measurement by the survey group of the sextupole
cores made by Process Equipment showed remarkably good adherence
to the specifications. The three quadrupole magnets for the energy-recovery
dump line (QIs) are all in final assembly at New England Technicoil
and are slated for delivery early next week.
In the corrector area, manufacture of the coils for the phasing
dipoles (DGs) started at MEI. However, manufacture of their mu
metal cores as well as the cores for the other reverse bend correctors
(DF&DCs) were rejected by the in house shop because of the
lack of large spot welding capability and tight tolerances. At
week's end we are going back to the mu metal supplier, negotiating
to utilize his spot welding and achievable tolerances to obtain
parts that will still meet our needs. Coils for the reverse bend
correctors were put on hold with the above problems but will now
be sent out for bid after a three week delay.
Drawings for the air core corrector coil sub assemblies were not
ready for signature as reported last week and are now in back
check. Manufacture of the prototype horizontal air core corrector
continued through the week, developing a method of winding the
coils in the flat and bending them around a mandrel in order to
achieve the saddle-like coil. At week's end it is clear that
this fabrication method requires too much development time to
achieve the required conductor placement and that direct three
dimensional winding, though requiring more tooling, will reliably
make the desired product. However, we will continue to make a
first prototype (that may not meet the required conductor placement)
with the flat wind method that will be adequate for a pivotal
cross talk test that duplicates the intended proximity to nearby
quadrupole's iron. Passing this test is needed before approval
of the final design. Design of the skew-quadrupole magnets remains
in final detailing.
Welding optical chicane chambers is complete and they are in cleaning.
Welding the bodies of the five straight reverse bend chamber
bodies is complete and their ends are being machined to accept
flanges. All parts for the curved reverse bend chambers are now
on hand and cleaned. Tacking of the bodies will start Monday.
Note that an incident of boiling from our vertical ultrasonic
cleaning station for vacuum tubes has shut down the facility,
snagging our cleaning routine for a few days. The cleaning routines
have now changed to the off-line clean room as well as to the
extensive use of the low-particulate steam jenny developed earlier.
Note that magnetic tests on two completed optical chicane chambers
in a DW dipole indicate a magnet field enhancement offset in the
order of 5 parts in ten thousand. It varies at the parts in 100,000
level for various excitations. This value will be added to the
excitation curves for the dipoles.
All remaining stands and girders for all sections of the machine
are now designed and in fabrication. The next major stand order,
due the first week in September, is the stands and girders for
In the power supply area, the vendor supplying the power supplies
for injection and extraction dipoles has completed tests and will
ship by air freight next week with routing through customs requiring
an additional week .
In the installation area, the beam line in the injector dump
region is made up, with an offset at one bellows still needing
resolution. The closure of the injection line to the X chamber
is expected early next week.
Regarding cryomodule fabrication
The linac cryomodule instrumentation was checked out this week
in preparation for next week's planned cooldown.
Regarding RF systems:
In the next two weeks, the elbow for the 500 kW High Voltage Power
Supply will be installed and tested in the unit. The Cathode
Power Supply for the buncher system is being wired for the required
PSS interlocks. The RF system for the buncher is about 50% complete.
The installation of the 50 kW klystron systems for the quarter
cryomodule are about 50% complete. We are still waiting for direction
from the vendor on what they want us to do with the second failed
50 kW klystron. AC power cables have been pulled but not connected
to the first Hipotronics 225 kW variable power supply. The required
cable lugs are being ordered. Hipotronics is sending two people
to make the necessary repairs to both units as a result of shipping
damage. The RF system for the linac is abut 40% complete. The
master oscillator is ready for phase noise testing with the drive
laser. Insulation of the LCW pipe and the M.O. distribution cables
are to be finished this month. The M.O. system is about 80% complete.
Regarding instrumentation and control systems:
The 1.5 micro Al foils are being mounted for the OTR beam viewers,
they look great! Foil this thin can be problematic, the mounting
fixtures are performing flawlessly. BPM's and beam viewers continue
to be mounted in the accelerator. The balance of the beam viewer
lenses have been ordered, these are a fixed focal length 100mm
The control and readout chassis for the 500 kV gun power supply
is in fabrication, this should be installed and tested by late
next week. This buffers the drive and readout analog signals
for the control system and provides a remote interface to the
control system. Check-out is complete for the cryomodule cooldown
diodes. A problem was found relating to an offset voltage, the
new VME crates do not have a ground connection to the power supplies,
the power supplies were referenced to ground and the problem was
solved. We are in contact with the manufacturer as to the reason,
other crates from the same shipment were found to have the same
The PSS is slated to be certified early next week, the wiring
and remote interfaces are being finished up this week.
The vacuum system control cables are now being pulled, the cable
tray has dividers mounted to segregate the high voltage cables
from the balance of the cables. A new cable tray was installed
this week to accommodate the power cables with the proper clearance
for cooling. All of the ion pump power supplies are mounted as
well as many of the vacuum control cards. The software controls
are in the works, total system testing will begin in a week or
The AMS backplane board is complete, signed off, and out for fabrication.
The pre-production AMS buffer board prototype was built in-house
and is being commissioned, the initial results demonstrate flat
response to 25 MHz.
The week was spent working on the drive laser and the optical system. The laser was leveled and locked into position, the pulse control and shaping system on the table is going together. The Installation Crew received and installed the stands for the drive laser transport. No interferences were encountered. We hung one mirror can to confirm we have the correct mounting hardware We will hang and position the two cans we have on hand next week. The last two cans are due next week, our Machine Shop confirms they will be on schedule. We received the complete order for the Brewster output windows for the User labs (six plus four spares). We also received part of the replacement order of metal mirrors that will be used in the mirror cassettes. We are setting up the downstream cavity assembly to check mirror positioning repeatability. We intend to move it into the FEL building as soon as the enclosures around the optical tables are installed. Michelle
Shinn gave a presentation on the FEL to staff at the Navy's Supervisor
of Shipbuilding in Newport News.
The installation and commissioning schedules were developed up
to the first light milestone. The slides are posted on the www
in the directory at http://www.jlab.org/~legg/fel_schedule.
The schedules were discussed in the commissioning meeting and
a suggestion was made that we test as much of the machine as possible
prior to the January scheduled down period to allow us to work
on problems during that period. It was agreed to examine this
approach for pitfalls during the coming week and to then make
The DW dipole problems were resolved to the satisfaction of the
measurement crew and the magnet performance integrator. The problem
was corrected by tightening the assembly bolts to reduce the magnet
to magnet gap variations. First article for installation on the
beamline should occur the first week in September.
We continued work with PARMELA to support the injector setup procedure.
The goal is to find the diagnostic which is most sensitive for
each parameter of the bunching/acceleration devices.
Punch list items continued to be worked. The ground rod conduit
water "leak" is expected to be corrected within the
next week. Telephones for the accelerator tunnel will be installed
on Friday. Elevator reinspection and completion date is currently
estimated to be September 5. Humidity control for the Clean Room
is being worked. The current problem involves a small (30 amp)
over-current protection device for the electric heat coils. Clean
Air Technologies has been contacted for some guidance. Resolution
is expected on Friday. Vinyl tile was installed in the elevator
landing at the accelerator level. Ceramic tile was installed in
the building foyer. The status of the civil construction effort
can be characterized as follows:
- The building is essentially complete including site paving
and final grading.
- All mechanical and electrical systems have been completed,
inspected and tested with the exception of the Direct Digital
Controls for the HVAC equipment and the emergency power generator.
- Punch lists items are being worked by the general contractor
and subs, however we continue to add items to the punch list.
Most punch list items are minor and will not be an issue to complete:
1) The perimeter drainage trench in the accelerator level does
not drain properly. A meeting has been arranged to resolve this
problem. 2) The concrete wall below the large glass curtain
wall adjacent to the front entrys was not installed vertical (the
concrete forms slipped during the placement of the concrete).